### v1

# Prompt-Guided Internal States for Hallucination Detection of Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains. However, they sometimes generate responses that are logically coherent but factually incorrect or misleading, which is known as LLM hallucinations. Data-driven supervised methods train hallucination detectors by leveraging the internal states of LLMs, but detectors trained on specific domains often struggle to generalize well to other domains. In this paper, we aim to enhance the cross-domain performance of supervised detectors with only in-domain data. We propose a novel framework, prompt-guided internal states for hallucination detection of LLMs, namely PRISM. By utilizing appropriate prompts to guide changes in the structure related to text truthfulness within the LLM’s internal states, we make this structure more salient and consistent across texts from different domains. We integrated our framework with existing hallucination detection methods and conducted experiments on datasets from different domains. The experimental results indicate that our framework significantly enhances the cross-domain generalization of existing hallucination detection methods [^1] We have open-sourced all the code, data, and models in GitHub: https://github.com/fujie-math/PRISM.
Prompt-Guided Internal States for Hallucination Detection of Large Language Models
1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains Dinan et al. (2018); Brown (2020); Zhang et al. (2022); Chowdhery et al. (2023). However, the hallucination problem in LLMs poses a potential threat to their practical application in many scenarios. LLM hallucinations refers to the cases where LLMs generate responses that are logically coherent but factually incorrect or misleading Zhou et al. (2020); Ji et al. (2023); Su et al. (2024). These hallucinated responses may be blindly accepted, leading users to learn incorrect information or take inappropriate actions. Therefore, detecting hallucinations in LLM-generated content becomes particularly meaningful. By effectively identifying incorrect generated content through a hallucination detector, users can be alerted to verify the accuracy of certain portions of the generated content and guide the LLM to regenerate relevant sections, thus preventing potential issues.
The unsupervised paradigm focuses on assessing the confidence of generated text and rejects the low-confidence ones. For example, the probability information for each token generated by LLMs can serve as a measure of hallucination Kadavath et al. (2022); Zhang et al. (2023); Quevedo et al. (2024); Hou et al. (2024). Additionally, the consistency among multiple responses generated by the LLM to the same question can also be used to evaluate their confidence Manakul et al. (2023). Chen et al. (2024) shifts the judgment of consistency among multiple responses to their corresponding internal vectors. However, these methods often struggle to achieve ideal detection accuracy or require a significant amount of additional response time Su et al. (2024).
For this reason, researchers have begun exploring the use of data-driven supervised methods for detection. These methods are generally based on the assumption that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns within their internal states. Marks and Tegmark (2023) reveal that true and false statements have distinguishable geometric structures within the LLM, allowing us to build classifiers by learning this structure. Marks and Tegmark (2023) and Sky et al. (2024) utilize linear structures to distinguish between different statements. Azaria and Mitchell (2023) train fully connected networks to act as hallucination detectors by using activation values within the LLM.
Supervised detectors trained on specific domains often struggle to achieve good generalization performance in other domains. For example, Bürger et al. (2024) found that the structural differences in the LLM’s internal states for affirmative sentences differ significantly from those for negated sentences when distinguishing between true and false statements. To address this issue, many studies focus on constructing diverse datasets or performing feature selection within the LLM to achieve better generalization performance Chen et al. (2023); Bürger et al. (2024); Liu et al. (2024). However, these methods require collecting additional training data from other domains, which is resource-intensive. In this paper, we aim to answer the following research question:
Can we enhance the cross-domain performance of supervised detectors without using datasets from other domains?
Driven by this research question, we propose a novel framework called PRISM, which stands for prompt-guided internal states for hallucination detection of LLMs. By utilizing appropriate prompts to guide changes in the structure related to text truthfulness within the LLM’s internal states, we make this structure more salient and consistent across texts from different domains. This enables detectors trained in one domain to perform well in others without additional data. Our approach is based on the insight that while the internal states of LLMs encode rich semantic information, they are primarily optimized during pre-training to predict the next token rather than for hallucination detection. As a result, the directly extracted internal states contain a lot of domain-specific information unrelated to truthfulness, leading to detectors that are specific to certain domains and unable to achieve good cross-domain generalization.
Due to LLM’s powerful ability to understand and follow instructions, we explore the use of prompt-guided methods to generate internal states more focused on truthfulness. We employed various methods to investigate the effect of prompts from different perspectives. The findings indicate that the introduction of appropriate prompts can significantly improve the salience of the structure related to text truthfulness in the LLM’s internal states and make this structure more consistent across different domain datasets. We also provide a simple and effective method to generate and select appropriate prompts for hallucination detection tasks. By combining the prompt template with the text to be evaluated and inputting them into the LLM, we obtain internal states that are better suited as features for hallucination detection tasks. Subsequently, we can integrate with existing hallucination detection methods to construct suitable detectors for classifying the text.
In summary, the contributions of this paper are as follows: (1) We conduct an in-depth investigation into how specific prompts influence the internal state structures of LLMs with regard to text truthfulness. (2) We propose PRISM, a novel framework that utilizes prompt-guided internal states to enhance hallucination detection in LLMs. (3) Successfully integrated PRISM with existing hallucination detection methodologies, demonstrating its ability to significantly enhance generalization performance across various domains.
Figure 1: Visualization of the LLaMA2-7B-Chat model’s internal states by using 2-dimensional PCA on the True-False dataset before and after the introduction of Prompt 1, where label 0 represents false statements and label 1 represents true statements. Additionally, a logistic regression model is fitted to distinguish between true and false statements, with the decision boundary shown as a black dashed line.
2 Datasets
In our research, to study the cross-domain hallucination detection problem, we utilized several publicly available datasets that encompass data from various domains. The first dataset we used is the True-False dataset Azaria and Mitchell (2023), which consists of six sub-datasets: "animals," "cities," "companies," "elements," "facts," and "inventions." These sub-datasets share similar textual structures but contain content on different topics. Each sub-dataset includes almost the same number of true and false statements, such as: ’Meta Platforms has headquarters in the United States’ and ’Silver is used in catalytic converters and some electronic components.’ More detailed information about this dataset can be found in Appendix A.
Another dataset we used is from Bürger et al. (2024) and consists of 24 sub-datasets with 6 different topics: "animal_class," "cities," "inventors," "element_symb," "facts," and "sp_en_trans," as well as 4 different grammatical structures: affirmative statements, negated statements, logical conjunctions, and logical disjunctions. Among the affirmative statements, "element_symb," "animal_class," "inventors," and "facts" are subsets of the True-False datasets, while the other sub-datasets were constructed based on the datasets used by Marks and Tegmark (2023) and Levinstein and Herrmann (2024).
Due to data imbalance across topics, we focus more on grammatical structure variations within this dataset. Affirmative statements are structured similarly to the examples in the True-False dataset, while negated statements are formed by negating the affirmative statements using the word "not." The sentences in logical conjunctions and logical disjunctions were constructed by sampling and combining sentences from the affirmative statements and then connecting them with "and" or "or." The sampling rules were carefully designed to ensure the overall dataset is balanced between true and false statements. For clarity, we refer to this dataset as LogicStruct throughout this paper and present additional information about it in Appendix A.
3 Prompt-Guided Internal States
Several previous studies have focused on leveraging the internal states of LLMs for hallucination detection Azaria and Mitchell (2023); Liu et al. (2024); Su et al. (2024). They assume that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns within their internal states. Therefore, these studies typically select contextualized embeddings corresponding to specific tokens in certain hidden layers of LLMs, and then use them as features to train hallucination detectors. However, these embeddings were originally intended to guide LLMs’ text generation and they encode various information of the related text. As a result, the information related to text truthfulness becomes intertwined with domain-specific details, making it difficult for the detector to identify a consistent structure related to text truthfulness in LLMs’ internal states.
Therefore, we hope to find a way to guide changes in the LLM’s internal states such that the structure related to text truthfulness becomes more salient and consistent across texts from different domains. This would help the detector learn this structure and enhance its generalization. To achieve this goal, we experimented with a simple prompt:
Prompt 1:
Here is a statement: [statement]
Is the above statement correct?
This prompt directly asks the LLM the truthfulness of a specific statement. When we input it into the LLM, the generation of new tokens will revolve around this question, which will be reflected in the LLM’s internal states. Although the prompt itself cannot be classified as true or false, it contains statements with different truthfulness, so this difference will be still reflected in the internal states of the LLM, resulting in a new structure related to text truthfulness.
In the following two subsections, we will use various methods to demonstrate the changes in the structure related to text truthfulness within the LLM’s internal states before and after the introduction of the prompt and explain how these changes will effect the hallucination detection task. Our analysis is conducted on the sub-datasets of the True-False dataset. The feature vectors used in our analysis are the embeddings corresponding to the last token in the last layer of LLaMA2-7B-Chat. More experimental results under other conditions are presented in Section 5.
AnimalsCitiesCompaniesElementsFactsInventionsAverage000.050.050.050.050.10.10.10.10.150.150.150.150.20.20.20.20.250.250.250.250.30.30.30.3Variance RatioBeforeAfter
Figure 2: Comparison of variance ratios before and after the introduction of Prompt 1 on the True-False dataset.
Table 1: Cosine similarity between truthfulness directions across different topics in the True-False dataset, before and after Prompt 1 addition.
Cosine Similarity Between Truthfulness Directions Cosine Similarity After Prompt 1 Addition
Datasets Animals Cities Comp. Elements Facts Invent. Animals Cities Comp. Elements Facts Invent.
Animals 1.0000 0.4368 0.4668 0.5498 0.6950 0.3786 1.0000 0.8037 0.7589 0.7284 0.8345 0.8333
Cities 0.4368 1.0000 0.4122 0.3300 0.4707 0.2520 0.8037 1.0000 0.8349 0.7253 0.7274 0.8540
Comp. 0.4668 0.4122 1.0000 0.4302 0.5691 0.4102 0.7589 0.8349 1.0000 0.7717 0.7255 0.8210
Elements 0.5498 0.3300 0.4302 1.0000 0.6258 0.3301 0.7284 0.7253 0.7717 1.0000 0.8175 0.8071
Facts 0.6950 0.4707 0.5691 0.6258 1.0000 0.4757 0.8345 0.7274 0.7255 0.8175 1.0000 0.7950
Invent. 0.3786 0.2520 0.4102 0.3301 0.4757 1.0000 0.8333 0.8540 0.8210 0.8071 0.7950 1.0000
Average 0.5878 0.4836 0.5481 0.5443 0.6394 0.4744 0.8265 0.8242 0.8187 0.8083 0.8167 0.8517
3.1 Effect of the Prompt on Structure Salience
In this subsection, we apply principal component analysis and variance analysis in specific directions to reveal that the introduction of the prompt can make the structure related to text truthfulness within the LLM’s internal states more salient.
Some previous studies Marks and Tegmark (2023); Bürger et al. (2024) have shown that true and false statements have distinguishable geometric structures within the LLM on certain datasets, and these structures can be directly observed after applying Principal Component Analysis to reduce the dimensionality of the relevant data. Therefore, we first apply PCA to observe the salience of the structure related to text truthfulness before and after the introduction of Prompt 1.
From Figure 1, we can clearly observe that in all six datasets, the introduction of the prompt enables the first two principal components to effectively separate the true and false statements. Since the first two principal components represent the directions with the largest variance among the embeddings, this indicates that the introduction of the prompt makes the structure related to text truthfulness in the embeddings more salient. However, this is just an intuitive but imprecise perspective, as we cannot be certain about the actual meaning represented by the first two principal components.
To accurately describe the differences in structure salience, we conduct a more detailed variance analysis. We believe that the proportion of variance along a particular direction relative to the total variance can reflect the salient of the structure in that direction. The larger this proportion, the more salient the relevant structure is compared to others. Therefore, to measure the salience of the structure related to text truthfulness, we need to identify a direction in the vector space that is associated with this structure.
A simple method is to subtract the average embedding of false statements from the average embedding of true statements within the dataset. We can refer to this direction as the truthfulness direction. This direction was used by Marks and Tegmark (2023) for classifying statements with different truthfulness. Its specific formula is as follows:
$\theta=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}v_{i}^{+}-\frac{1}{N^{-}}\sum_{i=1}^{N^{-}}v_{i}^{-},$ (1)
where $\theta$ represents the truthfulness direction in dataset $D$, $v_{i}^{+}$ and $v_{i}^{-}$ denote the embeddings corresponding to true and false statements in dataset $D$, respectively, and $N^{+}$ and $N^{-}$ represent their counts.
Then we can perform variance analysis along this direction. Let $v_{i}$ represents the embedding corresponding to a statement within dataset $D$, and let $N$ denote the total number of statements. By calculating
$\hat{v_{i}}=v_{i}-\frac{1}{N}\sum_{j=1}^{N}v_{j},$ (2)
we can use the de-centered vectors $\hat{v_{i}}$ to form the data matrix $X^{T}=(\hat{v_{1}},\hat{v_{2}},\dots,\hat{v_{N}})$. Next, we compute the covariance matrix corresponding to the data matrix:
$\Sigma=\frac{1}{N-1}X^{T}X.$ (3)
In this way, we obtain the total variance among all embeddings in dataset $D$:
$V_{T}=\text{Trace}(\Sigma)$ (4)
and the variance along the truth direction:
$V_{\theta}=\frac{\theta^{T}\Sigma\theta}{\|\theta\|^{2}}.$ (5)
Finally, we can use the ratio
$R=\frac{V_{\theta}}{V_{T}}$ (6)
to represent the salience of the structure related to text truthfulness. We calculate the corresponding variance ratios on the six sub-datasets of the True-False dataset and presented the results in Figure 2.
We can see that after the introduction of the prompt, this ratio shows a significant increase for all sub-datasets, indicating that the differences among the data begin to concentrate more along the truthfulness direction. This enhances the salience of the structure related to text truthfulness in these embeddings, making it easier for the hallucination detector to learn.
3.2 Effect of the Prompt on Structural Consistency
In this subsection, we focus on analyzing the consistency of structures related to text truthfulness across different datasets. We calculated the cosine similarity between truthfulness directions and observed that the introduction of the prompt makes these structures more consistent.
The salience of the structure related to text truthfulness may impact the training performance of the hallucination detector. However, when we aim to apply a detector trained in one domain to others, the consistency of this structure becomes particularly important. A more consistent structure will lead to better generalization performance. Therefore, we continue to use the ’truthfulness direction’ defined by Equation (1) to analyze the consistency of this structure across different datasets. We use cosine similarity to represent this consistency:
$c_{ij}=\frac{\theta_{i}\cdot\theta_{j}}{\|\theta_{i}\|\|\theta_{j}\|},$ (7)
where $c_{ij}$ represents the cosine of the angle between the truthfulness directions of datasets $D_{i}$ and $D_{j}$.
The calculation results for the six sub-datasets of the True-False dataset are presented in Table 1. We can see that the introduction of the prompt significantly increases the cosine similarity between the truthfulness directions across datasets, indicating that the truthfulness directions between different datasets become more consistent. This consistency will help improve the generalization of the related detector, which uses the LLM’s internal states as input features to determine the truthfulness of statements.
Table 2: Average variance ratios for 10 candidate prompt templates from Appendix C along with their ranking.
Prompts I II III IV V VI VII VIII IX X Without
Ratio 0.1093 0.1312 0.1291 0.1407 0.0875 0.1165 0.1041 0.1265 0.1228 0.1513 0.0725
Ranking 8 3 4 2 10 7 9 5 6 1 11
4 Methodology
The analyses in sections 3 indicate that the introduction of appropriate prompts can guide changes in the structure related to text truthfulness within the LLM’s internal states, and these changes will facilitate our use of LLMs’ internal states as features to perform the hallucination detection task. Based on this analysis, we propose a novel framework PRISM, i.e., PRompt-guided Internal States for hallucination detection of Large Language Models. This framework can be effectively integrated with existing hallucination detection methods to enhance their generalization across different domain data.
Our framework consists of two components. First, we manually construct a prompt template $P$ related to the hallucination detection task, and then we use the language model $L$ to generate multiple candidate prompt templates $P_{i}$ that are similar in meaning but different in structure. For each $P_{i}$, we combine it with text $a_{j}$ from the labeled dataset $A$ to obtain text $P_{i}(a_{j})$, which is then input into $L$ to obtain the corresponding feature vector $v_{ij}$. By using all feature vectors to calculate the structure salience index $R_{i}$ corresponding to each $P_{i}$ on $A$, we select the $P_{i}$ with the highest $R_{i}$ as our final prompt template $P_{s}$.
Next, we integrate the existing hallucination detection methods into our framework, training the hallucination detector $D$ on $A$ using all the feature vector $v_{j}$ and their label $b_{j}$ corresponding to the text $P_{s}(a_{j})$. For new text $T$ to be detected, we similarly combine it with the prompt template $P_{s}$ to obtain the text $P_{s}(T)$ and input it into the language model $L$ to obtain the corresponding feature vector $v$. Finally, we input $v$ into the hallucination detector $D$ to get its hallucination label $H$.
The above describes the complete process of our framework. Next, we will provide a more detailed explanation of some key steps in this framework.
4.1 Prompt Generation
We need to generate a large number of suitable candidate prompt templates for the hallucination detection task, and we can use LLMs to accomplish this process. First, we manually construct a simple prompt, such as Prompt 1 in Section 3. Then, we use this prompt to construct the following prompt:
Prompt 2:
"Here is a statement: ’[statement]’
Is the above statement correct?"
This is a universal prompt template. Please
generate templates with similar meanings but
diverse forms. The template should include the
embedding position of [statement].
Finally, we input the above prompt into the LLM to obtain a large number of prompt templates similar to Prompt 1. For example, we used Prompt 2 to query GPT-4o, resulting in 10 prompt templates that have similar meanings but different formats, with the relevant content presented in Appendix C.
Table 3: The experimental results on the True-False dataset using LLaMA2-7B-Chat.
Baselines Animals Cities Comp. Elements Facts Invent. Average
LN-PP 0.5496 0.5653 0.5357 0.5727 0.5948 0.5527 0.5618
MIND 0.4626 0.4664 0.4997 0.4850 0.5025 0.5023 0.4864
MM 0.5300 0.5688 0.5648 0.5581 0.6137 0.5075 0.5572
SAPLMA 0.6539 0.6700 0.6262 0.6057 0.7585 0.6233 0.6563
PRISM-MM 0.7004 0.9060 0.7908 0.6385 0.7031 0.7756 0.7524
PRISM-SAPLMA 0.7147 0.8936 0.8279 0.6705 0.7539 0.7804 0.7735
Table 4: The experimental results on the True-False dataset using LLaMA2-13B-Chat.
Baselines Animals Cities Comp. Elements Facts Invent. Average
LN-PP 0.5538 0.5564 0.5397 0.5800 0.5954 0.5612 0.5644
MIND 0.4957 0.4705 0.4372 0.5011 0.4986 0.4502 0.4756
MM 0.5330 0.5366 0.5438 0.5327 0.6196 0.5085 0.5457
SAPLMA 0.6584 0.7065 0.6721 0.6472 0.8054 0.6253 0.6858
PRISM-MM 0.7111 0.8837 0.8498 0.7092 0.7804 0.7993 0.7889
PRISM-SAPLMA 0.7405 0.8960 0.8435 0.6997 0.7918 0.8209 0.7987
4.2 Prompt Selection
Once we have a large number of suitable prompt templates, we need to select the most appropriate one among them to perform our hallucination detection task. A simple and effective method is to calculate the variance ratio corresponding to each prompt template on the labeled dataset using Equation (6) and then leverage this ratio to guide our prompt selection. A higher ratio indicates that the prompt is more likely to achieve better results.
It is worth noting that this process relies on the selection of feature vectors from the internal states of the LLM. We selected the context embedding corresponding to the last token in the final layer of LLaMA2-Chat-7B as the feature vector and calculated the corresponding variance ratio for each of the 10 prompt templates from Appendix C. We computed the variance ratio for each prompt template across the six sub-datasets of the True-False dataset, averaged the results, and ranked them, as shown in Table 2.
From Table 2, we can see that all the prompt templates increased the variance proportion along the truthfulness direction in the True-False dataset. Among them, the 10th prompt template achieved the highest variance ratio and it is as follows:
Prompt 3:
Does the statement ’[statement]’ accurately
reflect the truth?
Therefore, we will select this prompt template for our framework in the experiments presented in the next section.
Table 5: The overall experimental results on the LogicStruct dataset.
LLaMA2-7B-Chat LLaMA2-13B-Chat
Baselines Neg. Conj. Disj. Average Neg. Conj. Disj. Average
LN-PP 0.4108 0.5278 0.5743 0.5043 0.4263 0.5243 0.5863 0.5123
MIND 0.5219 0.4318 0.4969 0.4835 0.4849 0.4873 0.5038 0.4920
MM 0.4948 0.4937 0.5013 0.4966 0.5071 0.4895 0.5020 0.4995
SAPLMA 0.4864 0.5196 0.5030 0.5030 0.5155 0.6788 0.5186 0.5710
PRISM-MM 0.5548 0.8087 0.5580 0.6405 0.5260 0.8734 0.6207 0.6734
PRISM-SAPLMA 0.7345 0.7529 0.5329 0.6734 0.7389 0.8112 0.5722 0.7074
5 Experiment
5.1 Dataset and Metric
We evaluate our proposed framework PRISM and other baselines on the True-False and LogicStruct datasets which are detailed in section 2. Since the ratio of the two classes of labels in these datasets is close to 1:1, classification accuracy is an appropriate metric.
5.2 Baselines
We choose the following hallucination detection methods as baselines:
  • LN-PP Manakul et al. (2023) detects hallucinations based on the average probabilities of tokens generated by the LLM. We first determine a threshold, and then use this threshold for label classification.
  • SAPLMA Azaria and Mitchell (2023) is an effective supervised learning method. It uses activation values within the LLM as feature vectors to train a fully connected network as the hallucination detector. Selections of activation values from different layers can all achieve good classification performance.
  • MIND Su et al. (2024) is an unsupervised learning framework. It automatically generates labeled datasets from high-quality Wikipedia articles for training the hallucination detector. Other settings are consistent with the SAPLMA method.
  • MMMarks and Tegmark (2023) uses a simple mass-mean probe to distinguish between true and false statements. This method selects specific activation values as feature vectors, then uses Equation (1) to calculate the truthfulness direction on the labeled dataset. For the text to be detected, it first projects the feature vector onto the truthfulness direction by taking the dot product, then applies the sigmoid function, and finally rounds the result to obtain the label.
  • PRISM. Our framework selects Prompt 3 from Section 4.2 as the prompt template, and integrates with the SAPLMA and MM methods to obtain the PRISM-SAPLMA and PRISM-MM methods.
5.3 Implementation Details
We conducted our experiments using LLaMA2-7B-Chat and LLaMA2-13B-Chat Touvron et al. (2023). The activation values were selected from the context embeddings corresponding to the last token in the last layer of the LLMs. In the SAPLMA and MIND methods, the dimensions of the detectors for each layer were 256, 128, 64, and 2, with a dropout rate of 20% applied in the first layer. The activation function was ReLU, and the optimizer was Adam.
Our experiments focused on the generalization performance of different methods across different domains. For the SAPLMA method, we trained the detector separately on each subset of the True-False dataset and tested it on all other subsets. Thus, for a single subset, we obtained multiple test results from detectors trained on other subsets and averaged these values to obtain the final accuracy for this subset. In the LogicStruct dataset, we concentrated on training with affirmative statements and testing on other grammatical structures, as previous studies have indicated that achieving good generalization results in this scenario is challenging Marks and Tegmark (2023); Bürger et al. (2024). Both the MM and LN-PP methods employed the same evaluation approach as SAPLMA. The MIND method was trained on its automatically generated dataset and tested on each subset of the True-False and LogicStruct datasets. The settings of the methods under PRISM framework remain consistent with those of the original methods.
For all detectors that required training, we split a validation set from the training set at a 4:1 ratio and selected the model parameters with the highest accuracy on the validation set over 10 training epochs as the test parameters. The final results presented are the averages obtained from training under three different random seeds.
5.4 Experimental Results
We can see from Table 3 and 4 that our framework significantly outperforms other baselines in terms of average accuracy on the True-False dataset. On almost every sub-dataset of different topics, our framework substantially improves the generalization performance of the original methods. Additionally, there is a significant difference in the performance of the original MM and SAPLMA methods, which may be due to the simpler structure of the detector used in the MM method compared to that in the SAPLMA method. However, after integrating both into our framework, these two methods achieved similar performance. This indicates that the introduction of prompts made the structure related to text truthfulness more salient in the dataset, making it easier for detectors to capture. On the Facts dataset, the SAPLMA method and our framework achieve nearly identical accuracy. As shown in Figure 1, before the introduction of prompts, the true and false statements on the Facts dataset already exhibit clear separability, making it difficult to further enhance this structure.
Table 5 presents the experimental results on the LogicStruct dataset. We can observe that, except for our framework, other baselines struggle to generalize the training results on affirmative statements to other grammatical structures. According to the observations made by Bürger et al. (2024), this may be because the structure related to text truthfulness in affirmative statements is different from the corresponding structure in other grammatical structures. In contrast, our framework achieves significant classification performance across these grammatical structures, indicating that the introduction of prompts indeed makes the structure related to text truthfulness more consistent across different datasets. Additionally, we observe that the average results on the 13B model outperform those on the 7B model, suggesting that as the model’s reasoning capabilities improve, our framework is likely to perform better as well.
5.5 Ablation Studies
5.5.1 Impact of Prompt Selection
We conducted empirical experiments using the PRISM-SAPLMA method on all 10 prompt templates in Appendix C, with the results presented in Table 6. We can observe that all the prompt templates significantly improved the average accuracy compared to the SAPLMA method in Table 3, indicating that the improvement in the generalization performance of the hallucination detector due to the prompts is general and stable.
Additionally, by combining the ranking information from Table 2 and Table 6, we can see that the optimal prompt template selected based on the variance ratio also achieved the best results in the actual experiments. Furthermore, the 7th prompt template performs significantly worse in the actual experiments compared to others, while we ranked it 9th based on the variance ratio. These results indicate that the variance ratio can help us select prompt templates that achieve better generalization results in actual experiments while avoiding poorly performing ones.
5.5.2 Impact of Layer Selection
In addition to selecting context embeddings corresponding to the last token in the last layer of LLMs as feature vectors, we also extracted embeddings corresponding to the last token from the middle layer (16th) of LLaMA2-7B-Chat. We conducted experiments on the True-False dataset with our framework and the corresponding original methods, keeping all other settings the same, and the results are as follows:
Baselines Average
MM 0.5166
SAPLMA 0.5766
PRISM-MM 0.7924
PRISM-SAPLMA 0.7938
Compared to Table 3, the performance of the original methods has worsened, while our framework has achieved better results. This indicates that our framework exhibits good stability with respect to different layer selections.
Table 6: The experimental results of all 10 prompt templates in Appendix C on the True-False dataset using LLaMA2-7B-Chat.
Prompt Animals Cities Comp. Elements Facts Invent. Average Ranking
I 0.6889 0.8823 0.8302 0.7066 0.7475 0.7721 0.7713 2
II 0.6395 0.8915 0.8138 0.6318 0.7419 0.7183 0.7395 7
III 0.6460 0.9021 0.8087 0.6394 0.6956 0.7016 0.7322 8
IV 0.6887 0.8959 0.8058 0.6869 0.7436 0.7797 0.7668 3
V 0.7004 0.8506 0.8005 0.6639 0.7550 0.7123 0.7471 6
VI 0.6895 0.9115 0.7903 0.6699 0.7396 0.7587 0.7599 5
VII 0.6540 0.7764 0.7053 0.6200 0.7406 0.6532 0.6916 10
VIII 0.6829 0.9239 0.8205 0.6809 0.7318 0.7347 0.7624 4
IX 0.6310 0.8594 0.7979 0.6303 0.7031 0.7011 0.7205 9
X 0.7133 0.8930 0.8297 0.6736 0.7501 0.7801 0.7733 1
5.5.3 Impact of the Internal States
After introducing the prompt, is it possible to assess the truthfulness of the text directly based on the LLM’s response, without relying on its internal states? To investigate this question, we directly extract the token probabilities for " Yes" and " No" from LLaMA2-7B-Chat when generating the next token and compute their ratio $p[3869]/p[1939]$, where $p$ represents the vocabulary probabilities for generating the next token. We use whether this ratio is greater than or equal to 1 to determine the label of the original text. The average accuracy on the True-False dataset is as follows:
Baselines Average
Yes/No 0.6166
PRISM-MM 0.7524
PRISM-SAPLMA 0.7735
We can observe a significant accuracy difference between directly using the responses generated by the LLM and training the detector based on the LLM’s internal states. This indicates that the LLM’s internal states contain more information related to text truthfulness, which is not observable from the LLM’s responses but still plays a crucial role in hallucination detection.
6 Conclusions
In this paper, we first investigate the effect of specific prompts on the structure related to text truthfulness within the internal states of LLMs. We find that suitable prompts can make this structure more salient and consistent across datasets from different domains, facilitating hallucination detectors to learn this structure and improve generalization. Based on this insight, we introduce a novel framework, prompt-based internal states for hallucination detection of large language models, namely PRISM. This framework initially generates suitable prompt templates for hallucination detection task through a simple and effective approach, then integrates with existing hallucination detection methods to achieve detectors with improved generalization. We conducted experiments across various baselines on datasets from different domains, and the results demonstrate that our framework significantly enhances the generalization performance of existing methods and remains stable with different prompt and feature selections. We hope our framework can inspire future research in LLMs hallucination detection.
Limitations
We acknowledge that our research still has certain limitations. When detecting hallucinations in content generated by LLMs, we rely on the model’s internal states, which can be challenging to access in some cases and may require the use of a proxy model. Additionally, our approach does not leverage other information produced by the LLM during text generation, such as probability data and the generated text itself. In future research, we will explore fully utilizing various types of information produced by LLMs during text generation to achieve more effective hallucination detection.
Ethics Statement
The development and application of the PRISM framework for hallucination detection in large language models (LLMs) are guided by the principle of assessing the reliability and credibility of AI-generated content. PRISM aims to reduce the risk of users being misled by AI-generated outputs, thereby contributing to safer and more trustworthy AI applications. Additionally, the datasets used in this project are publicly available and do not involve personal or sensitive data, ensuring privacy and security. We need to recognize that our detector classifies based on the internal states of the model, and thus may inherit biases from the language model on certain issues, potentially leading to incorrect judgments by the detector. Therefore, we encourage further research and collaboration to develop ethical, fair, and trustworthy AI systems.
Appendix A Dataset Details
The following are some examples of statements from the True-False dataset:
  • Animals: "The gazelle has distinctive orange and black stripes and is an apex predator."
  • Cities: "Oranjestad is a city in Aruba."
  • Companies: "Meta Platforms has headquarters in United States."
  • Elements: "Silver is used in catalytic converters and some electronic components."
  • Facts: "The planet Jupiter has many moons."
  • Inventions: "Narinder Singh Kapany invented the programming language (theoretical)."
We can observe that the displayed statements contain both true and false information, and the relevant details are shown in Table 7.
Table 7: The information of the True-False dataset, where T=0 indicates that the sentence is annotated as false, and T=1 indicates true.
Topic #Sentence #T=0 #T=1
Animals 1008 504 504
Cities 1458 729 729
Companies 1200 600 600
Elements 930 465 465
Facts 613 307 306
Inventions 876 412 464
Examples of sentences with different grammatical structures in the LogicStruct dataset are as follows:
  • Affirmative statements: "The Spanish word ’dos’ means ’enemy’."
  • Negated statements: "The Spanish word ’dos’ does not mean ’enemy’."
  • Logical conjunctions: It is the case both that [statement 1] and that [statement 2].
  • Logical disjunctions: It is the case either that [statement 1] or that [statement 2].
In both logical conjunctions and logical disjunctions, Statement 1 and Statement 2 are derived from affirmative statements and sampled in a way that ensures label balance across the overall dataset. Additional information is presented in Table 8.
Table 8: The information of the LogicStruct dataset, where {Topic} in the Name refers to "animal_class," "cities," "inventors," "element_symb," "facts," and "sp_en_trans," and #Sentence represents the total number of data across all topics for each grammatical structure.
Name Description #Sentence
{Topic} Affirmative statements 3167
{Topic}_neg Negated statements 3167
{Topic}_conj Logical conjunctions 3998
{Topic}_disj Logical disjunctions 3000
Appendix B Analysis of Prompt 3
Consistent with the analysis of Prompt 1 in Section 3, we computed the ratio of the variance along the truthfulness direction to the total variance for the six sub-datasets of the True-False dataset. The results before and after the introduction of Prompt 3 are presented in Figure 3. We also computed the cosine similarity of the truthfulness direction between these datasets after the introduction of Prompt 3, as shown in Table 9.
AnimalsCitiesCompaniesElementsFactsInventionsAverage000.050.050.050.050.10.10.10.10.150.150.150.150.20.20.20.20.250.250.250.250.30.30.30.3Variance RatioBeforeAfter
Figure 3: Comparison of variance ratios before and after the introduction of Prompt 3 on the True-False dataset.
We can observe that Prompt 3 significantly enhances the salience and consistency of the structure related to text truthfulness in the embeddings. This is consistent with the analysis of Prompt 1 in Section 3, indicating that using prompts to guide changes in this structure within the LLM’s internal states is generally effective.
Table 9: Cosine similarity between truthfulness directions across different topics in the True-False dataset after Prompt 3 addition.
Datasets Animals Cities Comp. Elements Facts Invent.
Animals 1.0000 0.8370 0.8043 0.8144 0.8519 0.8811
Cities 0.8370 1.0000 0.8113 0.7835 0.8570 0.8613
Comp. 0.8043 0.8113 1.0000 0.8012 0.7927 0.8192
Elements 0.8144 0.7835 0.8012 1.0000 0.8327 0.8154
Facts 0.8519 0.8570 0.7927 0.8327 1.0000 0.8522
Invent. 0.8811 0.8613 0.8192 0.8154 0.8522 1.0000
Average 0.8265 0.8242 0.8187 0.8083 0.8167 0.8517
Appendix C Generation of Prompt Templates
Question: "Here is a statement: ’[statement]’ Is the above statement correct?" This is a universal prompt template. Please generate templates with similar meanings but diverse forms. The template should include the embedding position of [statement]. GPT-4o: Here are various templates with diverse forms that ask about the correctness or validity of a given statement: 1. "Is it accurate to say that ’[statement]’?" 2. "Would you consider the statement ’[statement]’ to be correct?" 3. "Can we confirm that ’[statement]’ is true?" 4. "Does the statement ’[statement]’ hold true?" 5. "Is ’[statement]’ a valid statement?" 6. "Is there accuracy in the claim ’[statement]’?" 7. "Could ’[statement]’ be considered a factual statement?" 8. "Is it correct to assume that ’[statement]’ is true?" 9. "Would it be right to say ’[statement]’ is accurate?" 10. "Does the statement ’[statement]’ accurately reflect the truth?" Each version maintains the focus on verifying the truth or accuracy of the statement in question.